OpenStudy (anonymous):

simplify x/x^2-25 + 5/x(x-5)

OpenStudy (anonymous):

\[ \frac{ X }{ \frac{ X ^{2}-25 }{ \frac{ 5 }{ X(X-5) } } }\]

OpenStudy (anonymous):

that ?

OpenStudy (anonymous):

yes but im sorry instead of the division its addition

OpenStudy (anonymous):

which division exactly ^_^

OpenStudy (anonymous):

the one on top of 5

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{ X ^{2}+5X+25 }{ X(X+5)(X-5) }\]